# Optical Image-to-Image Translation Using Denoising Diffusion Models: Heterogeneous Change Detection as a Use Case

### v1

@@ V-A1 Image Quality Metrics and Ablation Experiments
(b) Pre-Event Synthetic Planet Dove (Regression + $\mathbf{W}$)
@@
V-A2 Visual Examples
To properly convey the visual differences between the tested methods in Table II, Figure 4 shows generated examples of a region in Beirut, between (4b to 4f), and put them beside the corresponding Sentinel-II image (4a). As expected, regression-generated images (4b and 4c) have a blurry appearance and, due to this, when compared to the Sentinel-II version, an improvement in the apparent resolution is not easily noticeable. Figure 4b seems to have better inter-patch consistency, whereas the addition of whitening and coloring processes apparently brought some contrast enhancement in Figure 4c. Figure 4d is the result for the vanilla DDIM, where it can be noticed how the model has serious issues trying to figure out which is the overall tonality of the produced patches, thus creating a mosaic-like appearance when they are joined together to form a bigger image. This issue is solved by the inclusion of whitening and coloring processes, as observed by the result in Figure 4e. In fact, with their inclusion, 4e shows an even greater patch consistency than the regression-based results 4b and 4c. When compared to 4f, however, some regions of Figure 4e show deficiencies. In the center-west part of 4e, we observe a blurry and watched-out patch that apparently was not properly generated by the model. Likewise, in the south-west and south-east we observe two patches that are darker than their neighbors, and, for the south-east patch in particular, the darker patch also shows blurry low-resolution features that are not present in the same region in 4f. Overall, the PSNR voting procedure has improved patch consistency and suppressed non-ideal DDIM generations.
@@ V-A2 Visual Examples
(b) Pre-Event Synthetic Planet Dove (Regression + $\mathbf{W}$)
@@
Figures 6 and 7 respectively show images from the selected regions of Beirut and Austin. Their first, second, third and fourth rows individually show pre-event Sentinel-II, pre-event Synthetic Planet Dove (Regression + $\mathbf{W}$), pre-event Synthetic Planet Dove (Proposed), pre-event Planet Dove, post-event Planet Dove, and the change detection ground truth obtained by applying Algorithm 5, which is discussed in Section V-B, using Planet Dove pre and post event images. Instead of manual annotations, we use this ground truth since our I2I translation method aims to closely match low-resolution input to the style and resolution of high-resolution images, specifically Planet Dove in our experiments. In both regions, the tonality of the synthetic images generated using the proposed method (Figs. 6c and 7c) closely resemble those of the Planet pre-event images (Figs. 6d and 7d). Also, when they are compared with the pre-event Sentinel-II images used as inputs to the DDM-based method (Figs. 6a and 7a), it is easily observable how they have an improved apparent resolution and contrast. It is also noticeable how no perceptible undesired hallucinated features are present, favoring their use for change detection. Compared to the Regression + $\mathbf{W}$ synthetic images (Figs. 6b and 7b), our method delivered images with higher contrast and visual resolution, which translated into images that better match the characteristics of Planet Dove imagery.
@@
V-B Change Detection
V-B1 Procedure
@@
V-B2 Experiments
Two pairs of Sentinel-II pre-event and Planet Dove post-event images from two different cities were selected for a change detection comparison experiment. We also use their respective pre-event Planet Dove images to produce a change detection ground truth. The Sentinel-II pre-event images are part of the test set mentioned in Section IV. The first pair of images is in Beirut, Lebanon, surrounding the port region. It has been extracted from the northern part of the images in Figure 1, where no clouds are present. This image has been considered because it showcases multiple different types of changes, making targetless change detection useful. Moreover, it is a good example to test the robustness of the proposed method. The second pair of images is in Austin, USA, where many construction and terrain-related changes are present. The metrics used henceforth are the detection rate (DR) (recall) and the false alarm rate (FAR), which is defined as the ratio between the number of pixels misclassified as changes and the total number of unchanged pixels. These are metrics that properly quantify the trade-off between over and under detection.
@@ V-B2 Experiments
(b) Regression + $\mathbf{W}$ Synthetic Planet (Pre) + Planet (Post)
@@
Using the images presented in Figures 6 and 7, quantitative change detection results were obtained and presented in Figures 8 and 10 for Beirut and Austin, respectively. Pixels in green correspond to true positives, in red to false positives, and in blue to false negatives. The first, second, and third rows contain, individually, change detection results generated using as pre-event images: a Sentinel-II image (Figs. 8a and 10a), a synthetic Planet Dove image using the Regression + $\mathbf{W}$ model (Figs. 8b and 10b), described in the previous section, and a synthetic Planet Dove image generated with the proposed DDM-based algorithm (Figs. 8c and 10c). The Coloring procedure executed in the generation of the synthetic images made use of color information extracted from the post-Planet Dove image patches. A threshold $\omega$ has been chosen for each generated change map such that $\text{DR}=0.75$, which is an operating point that highlights the effects of the proposed algorithm.
@@
(a) Sentinel-II (Pre) + Planet (Post)
(b) Regression + $\mathbf{W}$ Synthetic Planet (Pre) + Planet (Post)
(c) Proposed Synthetic Planet (Pre) + Planet (Post)
(d) Ground Truth—Planet Pre and Post
To better evaluate the performance of the proposed algorithm for all possible choices of clipping threshold $\omega$, employed in Line 10 of Algorithm 5, Figures 9a and 9b present receiving operating characteristic (ROC) comparisons between the baseline change detection and change detection using the proposed method, for Beirut port and Austin regions, respectively. These curves were obtained by varying $\omega$ between 0 and 1.0, with steps of 0.01. As can be visualized in the curves, $\omega$ adjusts how permissive or restrictive should the algorithm be towards potential changes: a trade-off between higher detection rates and lower false alarm rates.
In Fig. 9a, the proposed method beats both compared methods for all choices of $\omega$, which displays how it has been able to significantly reduce false alarms for a desired detection rate. For fixed detection rates, the difference between the false alarm rate of the proposed method and the other two cases is up to 0.7 in the logarithmic scale or 5 times lower in the linear scale (DR=0.45).

### v2

@@ IV-A2 Inference Hyperparameters
(b) Pre-Event Synthetic Planet Dove (Regression + $\mathbf{W},\mathbf{C}$)
@@
IV-B Visual Examples
In our evaluation of image generation methods, as detailed in Table I, Figure 4 showcases the visual differences using examples from a region in Beirut. The sequence begins with our proposed method (4c), which utilizes the DDIM approach enhanced by whitening and coloring processes ($\mathbf{W},\mathbf{C}$) and PSNR voting, and ends with the SR3 diffusion-based method (4l). The Sentinel-II (4a) and Planet Dove (4b) images serve as input and target references, respectively.
Our method (4c) demonstrates superior patch consistency and detail resolution, outperforming the other techniques. The addition of $\mathbf{W},\mathbf{C}$ significantly enhances contrast and clarity, as evidenced by the comparison with the vanilla DDIM result (4e), which suffers from tonal inconsistencies and a mosaic-like appearance. These issues are notably absent in Figure 4d, where $\mathbf{W},\mathbf{C}$ have been applied.
Despite these improvements, some regions in Figure 4d exhibit deficiencies. The center-west patch is notably blurry and washed-out, suggesting a failure in the generation process. The south-west and south-east patches display a darker tone compared to adjacent areas, with the south-east patch also presenting blurry, low-resolution features not observed in the corresponding region of Figure 4c. This demonstrates how the PSNR voting mechanism improves overall generation quality.
The regression-based methods (4f and 4g), which share the same neural network backbone as the proposed method, yield images with a blurry appearance, which does not significantly enhance the apparent resolution when compared to the Sentinel-II image.
In a similar venue, the algorithms Pix2Pix (4h), ShuffleMixer (4i), SRDenseNet (4j), and SwinIR (4k) maintain good inter-patch consistency but fall short in capturing the high-resolution details that our method achieves. The SR3 method (4l) performs poorly, with color inconsistencies and feature representation problems that are more severe than those observed when our method is applied without $\mathbf{W},\mathbf{C}$.
@@ IV-B Visual Examples
(b) Pre-Event Synthetic Planet Dove (Regression + $\mathbf{W,C}$)
@@
Figures 6 and 7 respectively show images from the selected regions of Beirut and Austin. Their first, second, third and fourth rows individually show pre-event Sentinel-II, pre-event Synthetic Planet Dove (Regression + $\mathbf{W}$), pre-event Synthetic Planet Dove (Proposed), pre-event Planet Dove, post-event Planet Dove, and the change detection ground truth obtained by applying Algorithm 5, which is discussed in Section V, using Planet Dove pre and post event images. Instead of manual annotations, we use this ground truth since our I2I translation method aims to closely match low-resolution input to the style and resolution of high-resolution images, specifically Planet Dove in our experiments. In both regions, the tonality of the synthetic images generated using the proposed method (Figs. 6c and 7c) closely resemble those of the Planet pre-event images (Figs. 6d and 7d). Also, when they are compared with the pre-event Sentinel-II images used as inputs to the DDM-based method (Figs. 6a and 7a), it is easily observable how they have an improved apparent resolution and contrast. It is also noticeable how no perceptible undesired hallucinated features are present, favoring their use for change detection. Compared to the Regression + $\mathbf{W}$ synthetic images (Figs. 6b and 7b), our method delivered images with higher contrast and visual resolution, which translated into images that better match the characteristics of Planet Dove imagery.
Figure 8: Visualization of the intermediate steps of the diffusion process. The first row shows the progression of the noisy image $\bm{y}_{t}$ along diffusion timesteps. The second row shows their corresponding denoised predictions $\bm{y}_{0,t}$.
To demonstrate the dynamics of the forward and reverse diffusion processes for this specific task, Figure 8 shows an example of the intermediate steps of the denoising procedure. The first row shows the progression of the noisy image $\bm{y}_{t}$ along diffusion steps $t\in{0,1,...,T-1}$, whereas the second row shows their corresponding denoised outputs $\bm{y}_{0,t}$. During the very initial steps, most of the low-frequency information is recovered, resulting in blurry low-quality outputs. Around halfway of the process, the denoised image contains almost all high-frequency features of the final prediction. Thus, subsequent steps serve only to recover the finest details for the final prediction.
@@
V Change Detection as a Use Case
By transforming images from one domain to another, image-to-image translation methods can effectively highlight differences and changes between images taken at different times or under different conditions. Here, we explore the use of our image-to-image translation method for change detection, demonstrating its practical utility in this specific context. It is important to note that change detection is just one possible use case for our method. The results presented here aim to showcase how our image-to-image translation approach can be beneficial for practical applications, without limiting its potential to other domains.
V-A Rationale for Selective Comparisons
We focus on evaluating our model’s performance against one of the regression models previously compared: Regression + $\mathbf{W,C}$. Previous experiments have already demonstrated the superiority of our model in image generation compared to state-of-the-art algorithms. Specifically, our method outperforms others in image quality, particularly with respect to the LPIPS metric, which is a patch-by-patch comparative metric. A lower LPIPS score indicates better translation and fewer artifacts, which correlates with higher quality change detection results. To maintain a concise and focused manuscript, we limit our change detection experiments to comparisons against regression models. This approach allows us to highlight the power of using diffusion models against a regression-based objective, which is also employed by the methods ShuffleMixer, SRDenseNet, and SwinIR.
It is important to note that the tested change detection algorithm does not necessarily extract the full potential of our image translation method, as it contains steps that do not fully utilize the added visual resolution to our advantage. Additionally, using this specific simple HCD method with images generated by the proposed method does not necessarily produce better HCD results when compared to other methods such as ShuffleMixer, SRDenseNet, and SwinIR. These methods have performed poorly in super resolution but very well in patch-wise feature consistency, and Algorithm 5 does not make use of the higher resolution features produced by our image translation method. However, employing a more advanced change detection algorithm is beyond the scope of this paper. The HCD experiments are intended to demonstrate that our proposed method could be potentially useful for change detection. A more advanced learning-based change detection procedure using the produced synthetic images will be explored in future work. Furthermore, Regression + + $\mathbf{W,C}$ is trained with the same neural network backbone as the proposed method. By doing so, we effectively showcase the advantages of our diffusion-based method under consistent conditions, ensuring a clear and concise presentation of our results.
V-B Procedure
@@
(b) Regression + $\mathbf{W,C}$ Synthetic Planet (Pre) + Planet (Post)
@@
V-C Experiments
Two pairs of Sentinel-II pre-event and Planet Dove post-event images from two different cities were selected for a change detection comparison experiment. We also use their respective pre-event Planet Dove images to produce a change detection ground truth. The Sentinel-II pre-event images are part of the test set mentioned in Section III. The first pair of images is in Beirut, Lebanon, surrounding the port region. It has been extracted from the northern part of the images in Figure 1, where no clouds are present. This image has been considered because it showcases multiple different types of changes, making targetless change detection useful. Moreover, it is a good example to test the robustness of the proposed method. The second pair of images is in Austin, USA, where many construction and terrain-related changes are present. The metrics used henceforth are the detection rate (DR) (recall) and the false alarm rate (FAR), which is defined as the ratio between the number of pixels misclassified as changes and the total number of unchanged pixels. These are metrics that properly quantify the trade-off between over and under detection.
V-C1 Visual Evaluation
Using the images presented in Figures 6 and 7, quantitative change detection results were obtained and presented in Figures 9 and 10 for Beirut and Austin, respectively. Pixels in green correspond to true positives, in red to false positives, and in blue to false negatives. The first, second, and third rows individually contain change detection results generated using pre-event images: a Sentinel-II image (Figs. 9a and 10a), a synthetic Planet Dove image using the Regression + $\mathbf{W},\mathbf{C}$ model (Figs. 9b and 10b), described in the previous section, and a synthetic Planet Dove image generated with the proposed DDM-based algorithm (Figs. 9c and 10c). The Coloring procedure executed in the generation of the synthetic images made use of color information extracted from the post-Planet Dove image patches. A threshold $\omega$ has been chosen for each generated change map such that $\text{DR}=0.75$, which is an operating point that highlights the effects of the proposed algorithm.
(a) Sentinel-II (Pre) + Planet (Post)
(b) Regression + $\mathbf{W,C}$ Synthetic Planet (Pre) + Planet (Post)
(c) Proposed Synthetic Planet (Pre) + Planet (Post)
(d) Ground Truth—Planet Pre and Post
@@ V-C1 Visual Evaluation
V-C2 Quantitative Evaluation
To better evaluate the performance of the proposed algorithm for all possible choices of clipping threshold $\omega$, employed in Line 10 of Algorithm 5, Figures 11a and 11b present receiving operating characteristic (ROC) comparisons between the baseline change detection and change detection using the proposed method, for Beirut port and Austin regions, respectively. These curves were obtained by varying $\omega$ between 0 and 1.0, with steps of 0.01. As can be visualized in the curves, $\omega$ adjusts how permissive or restrictive should the algorithm be towards potential changes: a trade-off between higher detection rates and lower false alarm rates.
In Fig. 11a, the proposed method beats both compared methods for all choices of $\omega$, which displays how it has been able to significantly reduce false alarms for a desired detection rate. For fixed detection rates, the difference between the false alarm rate of the proposed method and the other two cases is up to 0.7 on the logarithmic scale or 5 times lower on the linear scale (DR=0.45).